gary is 5 years older than henry.21 years ago, gary was twice as old as henry.what is the age of gary and henry.

Answer by mananth(16946)

Let Henry be x
Gary = x+5
..
21 years ago their ages were
Henry x-21
Gary (x+5-21)
..
(x+5-21)= 2(x-21)
x-16 = 2x-42
add +16 to both sides
x-16+16=2x-42+16
x=2x-26
add -x to both sides
x-x=2x-x-26
0=x-26
add 26 to both sides
26=x-26+26
x= 26 Henry,s age
Gary = 26+5 = 31 years